# Stability of vehicular admission control schemes in urban traffic networks under modelling uncertainty

### v1

@@ VII Conclusions
This work considered a broad class of decentralized vehicular admission control schemes (VAC) described by general nonlinear dynamics and examined the stability properties of their feedback connection with large-scale, heterogeneous regional, traffic networks described by uncertain, nonlinear, concave macroscopic fundamental diagrams. Passivity theory was employed and scalable, locally verifiable conditions on the VAC dynamics were developed, that enable stability guarantees at the presence of modelling uncertainty. Suitable examples of VAC dynamics that demonstrate the applicability of the proposed conditions were presented. These conditions provide a flexible framework for the design of VAC schemes, enabling heterogeneous VAC dynamics in each region, and thus complementing different demand operational objectives, and network infrastructure. The developed analytic results were validated with numerical simulations on a 6-region system demonstrating the effectiveness and applicability of the presented VAC framework.
@@ Appendix C Proof of 2
The condition $\gamma^{c}_{i}<\beta^{c}_{i}/K_{i}$ in (31) implies that the $\mathcal{L}_{2}$-gain from $(\rho_{i}-\rho^{*}_{i})$ to $(p^{c}_{i}(\rho_{i})-p^{c}_{i}(\rho^{*}_{i}))$ is less than $\beta^{c}_{i}/K_{i}$ for any equilibrium $(\rho^{*}_{i},p^{c}_{i}(\rho^{*}_{i}))$. Hence, since (33) has an $\mathcal{L}_{2}$-gain of $K_{i}$, it follows that the $\mathcal{L}_{2}$ gain from $(\rho_{i}-\rho^{*}_{i})$ to $(u_{1,i}-u^{*}_{1,i})$ is less than $\beta^{c}_{i}$ for any feasible equilibrium pair $(\rho^{*}_{i},u^{*}_{1,i})$. This allows to show that the system from $\rho_{i}$ to $u_{i}$ is input strictly passive about any equilibrium by directly applying [KASIS7776980, Prop. 1]. ∎

### v2

@@ VII Conclusions
This work considered a broad class of decentralized vehicular admission control schemes (VAC) described by general nonlinear dynamics and examined the stability properties of their feedback connection with large-scale, heterogeneous regional, traffic networks described by uncertain, nonlinear, concave macroscopic fundamental diagrams. Passivity theory was employed and scalable, locally verifiable conditions on the VAC dynamics were developed, that enable stability guarantees at the presence of modelling uncertainty. Suitable examples of VAC dynamics that demonstrate the applicability of the proposed conditions were presented. These conditions provide a flexible framework for the design of VAC schemes, enabling heterogeneous VAC dynamics in each region, and thus complementing different demand operational objectives, and network infrastructure. The developed analytic results were validated with numerical simulations on a 6 and a 20-region system demonstrating the effectiveness and applicability of the presented VAC framework.
@@ Appendix C Proof of 2
The condition $\gamma^{c}_{i}<\beta^{c}_{i}/K_{i}$ in (31) implies that the $\mathcal{L}_{2}$-gain from $(\rho_{i}-\rho^{*}_{i})$ to $(p^{c}_{i}(\rho_{i})-p^{c}_{i}(\rho^{*}_{i}))$ is less than $\beta^{c}_{i}/K_{i}$ for any equilibrium $(\rho^{*}_{i},p^{c}_{i}(\rho^{*}_{i}))$. Hence, since (33) has an $\mathcal{L}_{2}$-gain of $K_{i}$, it follows that the $\mathcal{L}_{2}$ gain from $(\rho_{i}-\rho^{*}_{i})$ to $(u_{1,i}-u^{*}_{1,i})$ is less than $\beta^{c}_{i}$ for any feasible equilibrium pair $(\rho^{*}_{i},u^{*}_{1,i})$. This allows to show that the system from $\rho_{i}$ to $u_{i}$ is input strictly passive about any equilibrium by directly applying [22, Prop. 1]. ∎